OpenStudy (anonymous):

Write the equation in slope-intercept form: Write it leaving no spaces between the characters. 2x + 5y= 15 Answer

OpenStudy (blurbendy):

slope intercept form is y = mx + b 5y = -2x + 15 y = -2/5x + 3